OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 4x - 4y + 7 = 15 8y = 2x + 14

OpenStudy (anonymous):

do you know how to solve an equation with one variable ?

OpenStudy (anonymous):

are you trying to learn how to solve, or just want the answer ?

OpenStudy (anonymous):

Solve the first equation for x: 4x + 7 = 15 + 4y 4x = 8 +4y x = 2 + y Substituting this for x in the second equation gives: 8y = 2(2+y) + 14 distributive property: 8y = 4 + 2y + 14 8y = 2y + 18 6y = 18 y = 18/6 = 3 So we know that y = 3 Substituting for y in the first equation gives: 4x -4(3) +7 = 15 4x - 12 + 7 = 15 4x -5 = 15 4x = 20 x = 5